Chuck has $2,500 invested in a bank that pays 4% annually. How long will it take for his funds to double?

Answer by ewatrrr(24785)


Hi
A= p(1+r) ^t
A/p = 2 = (1.04)^t |money doubling
2 = (1.04)^t
ln2 = ln(1.04)^t |taking logarithm of each side of the equation
ln2 = t*ln(1.04)
ln2/ln(1.04) = t
17.673 yr = t, the time to double